# Learning a distance measure from the information-estimation geometry of data

### v1

@@ 4 Discussion
We have introduced the Information-Estimation Metric (IEM), a novel form of distance measure induced by the geometry of an underlying probability density, and provided a means of learning this (unsupervised) metric from samples. The definition of the IEM relies on fundamental principles that link the probability density with its local geometry, namely the pointwise I-MMSE and Tweedie–Miyasawa formulas. This relationship between probability and geometry arises from the choice of an estimation problem, in our case Gaussian denoising. Different choices of the estimation problem may yield different types of IEMs. For instance, it is possible to define an IEM using the pointwise I-MMSE relation for Poisson channels (Jiao et al., 2013), and the empirical Bayes relation for Poisson denoising (Raphan & Simoncelli, 2011). We leave these as opportunities for future work. Furthermore, the IEM does not assume that the density is supported on a low-dimensional manifold, in contrast to manifold learning approaches (see Sec. A.1 for further discussion). In fact, the IEM is well-defined for any valid probability density. We proved that the IEM is a valid distance metric and analyzed its local and global properties through both theoretical results and illustrative examples. To demonstrate the value of our proposed framework, we trained an IEM on the ImageNet database and found that it aligns surprisingly well with human judgments of image similarity.
@@
The proposed IEMs (Defs. 1 and 2) require choosing a scalar hyperparameter $\Gamma$. This hyperparameter sets the maximum SNR over which the integral is computed, effectively controlling the finest resolution at which the metric is adapted to the density. Such a hyperparameter should presumably be chosen based on the fine-scale geometry of the density, or for a learned density, the complexity and size of the training set. Moreover, the generalized IEM (Def. 2) depends on the choice of the function $f$, which qualitatively controls the relative importance of log-probability ratio values compared to score differences. A systematic principle for determining both $\Gamma$} and $f$ remains an open problem. Perhaps the most important limitation of the IEM is its computational cost: Numerical estimation of the integral is more computationally demanding than evaluation of existing supervised perceptual metrics (e.g., LPIPS or DISTS). This is acceptable for applications to collections of images (as in our comparison to human perceptual data), but would limit its use as an optimization objective (e.g., for solving inverse problems, or optimizing compression systems). We believe that evaluation of the IEM in a single forward pass may be possible—for example, using a strategy similar to Guth et al. (2025).
@@
There are many potential future applications for the IEM framework. For example, it offers new opportunities for unsupervised data clustering (as we illustrate in Sec. D.3), information retrieval, and for evaluating (or optimizing) image restoration and compression engines.
@@
Sections 3 and E provide all the details necessary to reproduce our results, including the training hyperparameters of the denoiser model used in the computation of the IEM, the implementation details of the learned function $f_{\omega}$, and the data preprocessing procedures. Our code is available online at https://github.com/ohayonguy/information-estimation-metric.
@@ A.1 Related work
Another path to metric learning is through dimensionality reduction, where the metric is derived by measuring distances (e.g., Euclidean) in a low-dimensional embedding space. For example, diffusion maps (Coifman & Lafon, 2006) aim to reveal the manifold structure of data. Given a similarity matrix and its corresponding graph, this method characterizes the geometry of the underlying manifold by simulating random walks. Computing the eigenvectors of the corresponding diffusion operator provides coordinates for embedding data into a lower-dimensional space, where Euclidean distances correspond to distances along the manifold and are referred to as “diffusion distances.” However, constructing the similarity matrix requires the choice of a kernel in data space, and the diffusion “time” needs to be carefullly calibrated (Shan & Daubechies, 2022). In contrast, the IEM does not make a manifold assumption and instead relies on denoising, which corresponds to reversing a diffusion process in the full signal space.
@@
Recent studies proposed different ways to define a Riemannian metric using the score of the blurred density (i.e., the density of noise-corrupted signals) (Saito & Matsubara, 2025; Azeglio & Bernardo, 2025). These Riemannian metrics are then used to compute geodesics, which can, e.g., be applied to interpolate between a given pair of images. Such approaches differ from our proposed IEM framework in several important ways. First, the IEM is, by definition, a global distance function from which we derive a local Riemannian metric, rather than the other way around. In fact, the IEM is not the geodesic distance corresponding to the associated Riemannian metric we derive (Eq. 5). Second, the Riemannian metrics proposed in (Saito & Matsubara, 2025; Azeglio & Bernardo, 2025) are defined based on the score at a single blur level $\gamma$, whereas the IEM (and its associated Riemannian metric) integrates over a range of blur levels $\gamma\in[0,{\Gamma}]$. Third, the IEM is constructed from first principles and satisfies several important properties, e.g., it reduces to the well-known Mahalanobis distance when the prior density is Gaussian.
@@
Information geometry (Amari & Nagaoka, 2000) uses tools from differential geometry to analyze statistical models. In particular, it considers probability distributions as points lying on a Riemannian manifold whose metric is derived from the KL divergence. In the case of a family of conditional distributions $\{p_{{\mathbf{y}}|{\mathbf{x}}}(\cdot|{\bm{x}})\,|\,{\bm{x}}\in\mathbb{R}^{d}\}$, this metric is given by the Fisher information matrix (Fisher, 1922), which quantifies the amount of information that $\mathbf{y}$} carries about $\mathbf{x}$} (here, $\mathbf{x}$} is considered as an unknown “parameter”). Specifically, the Fisher information matrix is defined as
@@
Solving a Bayesian inference problem requires both a likelihood function $p_{{\mathbf{y}}|{\mathbf{x}}}$ and a prior $p_{{\mathbf{x}}}$. However, in some cases only the likelihood is available. In such situations, it is natural to choose a non-informative prior using Jeffreys rule (Jeffreys, 1946), which states that the prior should be proportional to the square root of the determinant of the Fisher information matrix. More relevant to our case, this relationship has also been applied in the reverse direction, where the prior is known but the likelihood is not. In particular, it has been shown that a likelihood for which the Jeffreys prior matches the data distribution satisfies the principles of efficient coding (Ganguli & Simoncelli, 2014). The reverse use of Jeffreys prior has also been explored in machine learning. For example, Lebanon (2002) considered a Riemannian metric under which the data is uniformly distributed. Since the prior density is a scalar function, they assumed an isotropic metric of the form $\smash{{\bm{M}}({\bm{x}})\propto\lambda({\bm{x}}){\bm{I}}}$, where $\smash{\lambda({\bm{x}})\propto p({\bm{x}})^{2/d}}$, and ${\bm{x}}\in\mathbb{R}^{d}$. In both of these cases, the result depends only on a scalar quantity and therefore cannot account for the varying magnitudes of discrimination thresholds in different perturbation directions (Berardino et al., 2017). In contrast, the IEM builds on the local geometry of the data distribution to define a distance that captures its anisotropic structure.
@@ A.2 Origins of the pointwise I-MMSE formula
This reformulation highlights that the two sides of the I-MMSE formula correspond to random variables that are equal in expectation. Venkat & Weissman (2012) showed that these random variables satisfy the pointwise I-MMSE formula
@@
where this equality holds with probability one (i.e., almost surely). As noted by Venkat & Weissman (2012), taking expectations in Eq. 2 immediately recovers the original I–MMSE formula. Indeed, the stochastic integral on the right-hand side of Eq. 2 is a martingale with zero mean, while the left-hand side corresponds to the pointwise mutual information between $\mathbf{x}$} and ${\mathbf{y}}_{\gamma}$, whose expectation yields the mutual information between these two random vectors. Using the fact that
@@ Appendix B Mismatched IEM and local decomposition of Kullback–Leibler divergence
The mismatched IEM distance is directly related to the KL divergence between $p_{{\mathbf{x}}_{1}}$ and $p_{{\mathbf{x}}_{2}}$. Indeed, Venkat & Weissman (2012) showed that this KL divergence can be expressed as the expected quadratic variation of the log-probability ratio when the two distributions are evaluated at the same point, $\log\mathopen{}\mathclose{{\left(p_{{\mathbf{y}}_{1,\gamma}}(\gamma{\mathbf{x}}_{1}+{\mathbf{w}}_{\gamma})/p_{{\mathbf{y}}_{2,\gamma}}(\gamma{\mathbf{x}}_{1}+{\mathbf{w}}_{\gamma})}}\right)$, yielding
@@ C.2.2 From the Hessian of the noisy channel log-density to the posterior covariance
This relationship has already appeared in the literature in several contexts, and is often referred to as the “second-order Tweedie identity.” To the best of our knowledge, it was first derived by Hatsell & Nolte (1971, Proposition 3). For completeness and notational consistency, we include a derivation here.
@@ C.2.3 Average local metric
${\bm{e}}_{\gamma}({\bm{x}},{\mathbf{w}}_{\gamma})\coloneqq{\bm{x}}-\mathbb{E}[{\mathbf{x}}|{\mathbf{y}}_{\gamma}=\gamma{\bm{x}}+{\mathbf{w}}_{\gamma}]=-\frac{1}{\gamma}{\mathbf{w}}_{\gamma}-\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma}),$ (45)
@@
$-{\bm{e}}_{\gamma}({\bm{x}}+\bm{\epsilon},{\mathbf{w}}_{\gamma})$ $=\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})+\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\bm{\epsilon}$
$\quad{}+\frac{\gamma^{2}}{2}\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})(\bm{\epsilon},\bm{\epsilon})+o(\mathopen{}\mathclose{{\left\lVert\bm{\epsilon}}}\right\rVert^{2}),$ (46)
@@
$=\frac{1}{2}\int_{0}^{\infty}\mathbb{E}\mathopen{}\mathclose{{\left[2\mathopen{}\mathclose{{\left\langle\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma}),\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\bm{\epsilon}+\frac{\gamma^{2}}{2}\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})(\bm{\epsilon},\bm{\epsilon})}}\right\rangle}}\right$.
$\quad\quad\quad\quad\quad\quad\mathopen{}\mathclose{{\left.+\mathopen{}\mathclose{{\left\lVert\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\bm{\epsilon}}}\right\rVert^{2}}}\right]\mathrm{d}\gamma+o(\mathopen{}\mathclose{{\left\lVert\varepsilon}}\right\rVert^{2})$ (47)
$=\bm{\epsilon}^{\top}\int_{0}^{\infty}\mathbb{E}\mathopen{}\mathclose{{\left[\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)}}\right]\gamma\mathrm{d}\gamma$
$\quad+\frac{1}{2}\bm{\epsilon}^{\top}\mathopen{}\mathclose{{\left(\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma}}\right)\bm{\epsilon}$ (48)
@@
$-\nabla^{2}\log{p_{\mathbf{x}}({\bm{x}})}$ $=\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma.$ (49)
We then take expectation w.r.t. $p_{\mathbf{x}}$ and obtain
$\mathbb{E}\mathopen{}\mathclose{{\left[-\nabla^{2}\log p_{\mathbf{x}}({\bm{x}})}}\right]$ $=\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma.$ (50)
The second term in the expectation then vanishes, using Stein’s lemma (Stein, 1981),
@@
The last equality in the proposition is generic and classical. By expanding derivatives, we have
$-\int p_{{\mathbf{x}}}({\bm{x}})\nabla^{2}\log p_{{\mathbf{x}}}({\bm{x}})\mathrm{d}{\bm{x}}$ $=\int\mathopen{}\mathclose{{\left(p_{{\mathbf{x}}}({\bm{x}})\nabla\log p_{\mathbf{x}}({\bm{x}})\nabla\log p_{\mathbf{x}}({\bm{x}})^{\top}-\nabla^{2}p_{\mathbf{x}}({\bm{x}})}}\right)\mathrm{d}{\bm{x}},$ (54)
and the second term in the integral on the right-hand side vanishes through integration by parts.

### v2

@@ 4 Discussion
We have introduced the Information-Estimation Metric (IEM), a novel form of distance measure induced by the geometry of an underlying probability density, and provided a means of learning this (unsupervised) metric from samples. The definition of the IEM relies on fundamental principles that link the probability density with its local geometry, namely the pointwise I-MMSE and Tweedie–Miyasawa formulas. This relationship between probability and geometry arises from the choice of an estimation problem, in our case Gaussian denoising. Different choices of the estimation problem may yield different types of IEMs. For instance, it is possible to define an IEM using the pointwise I-MMSE relation for Poisson channels (Jiao et al., 2013), and the empirical Bayes relation for Poisson denoising (Raphan and Simoncelli, 2011). We leave these as opportunities for future work. Furthermore, the IEM does not assume that the density is supported on a low-dimensional manifold, in contrast to manifold learning approaches (see Sec. A.1 for further discussion). In fact, the IEM is well-defined for any valid probability density. We proved that the IEM is a valid distance metric and analyzed its local and global properties through both theoretical results and illustrative examples. To demonstrate the value of our proposed framework, we trained an IEM on the ImageNet database and found that it aligns surprisingly well with human judgments of image similarity.
@@
The proposed IEMs (Defs. 1 and 2) require choosing a scalar hyperparameter $\Gamma$. This hyperparameter sets the maximum SNR over which the integral is computed, effectively controlling the finest resolution at which the metric is adapted to the density. Such a hyperparameter should presumably be chosen based on the fine-scale geometry of the density, or for a learned density, the complexity and size of the training set. Moreover, the generalized IEM (Def. 2) depends on the choice of the function $f$, which qualitatively controls the relative importance of log-probability ratio values compared to score differences. A systematic principle for determining both $\Gamma$} and $f$ remains an open problem. Perhaps the most important limitation of the IEM is its computational cost: Numerical estimation of the integral is more computationally demanding than evaluation of existing supervised perceptual metrics (e.g., LPIPS or DISTS). This is acceptable for applications to collections of images (as in our comparison to human perceptual data), but would limit its use as an optimization objective (e.g., for solving inverse problems, or optimizing compression systems). We believe that the IEM may be evaluated in a single forward pass, e.g., using a strategy similar to Guth et al. (2025).
@@
There are many potential future applications for the IEM framework. For example, it offers new opportunities for unsupervised data clustering (as illustrated in Sec. D.4), information retrieval, evaluating (or optimizing) image restoration and compression engines, and discriminating between generative models (using the mismatched IEM proposed in App. B). It is also natural to consider applying the principles presented in this paper to other forms of continuous signals, such as audio.
@@
Sections 3, E and D provide all the details necessary to reproduce our results, including the training hyperparameters of the denoiser model used in the computation of the IEM, the implementation details of the learned function $f_{\omega}$, the data preprocessing procedures, and the maximum differentiation competition experiments. Our code is available online at https://github.com/ohayonguy/information-estimation-metric.
@@ A.1 Related work
Another path to metric learning is through dimensionality reduction, where the metric is derived by measuring distances (e.g., Euclidean) in a low-dimensional embedding space. For example, diffusion maps (Coifman and Lafon, 2006) aim to reveal the manifold structure of data. Given a similarity matrix and its corresponding graph, this method characterizes the geometry of the underlying manifold by simulating random walks. Computing the eigenvectors of the corresponding diffusion operator provides coordinates for embedding data into a lower-dimensional space, where Euclidean distances correspond to distances along the manifold and are referred to as “diffusion distances.” However, constructing the similarity matrix requires the choice of a kernel in data space, and the diffusion “time” needs to be carefully calibrated (Shan and Daubechies, 2022). In contrast, the IEM does not make a manifold assumption and instead relies on denoising, which corresponds to reversing a diffusion process in the full signal space.
@@
Recent studies proposed different ways to define a Riemannian metric using the score of the blurred density (i.e., the density of noise-corrupted signals) (Saito and Matsubara, 2025; Azeglio and Bernardo, 2025). These Riemannian metrics are then used to compute geodesics, which can, e.g., be applied to interpolate between a given pair of images. Such approaches differ from our proposed IEM framework in several important ways. First, the IEM is, by definition, a global distance function from which we derive a local Riemannian metric, rather than the other way around. In fact, the IEM is not the geodesic distance corresponding to the associated Riemannian metric we derive (Eq. 5). Second, the Riemannian metrics proposed in (Saito and Matsubara, 2025; Azeglio and Bernardo, 2025) are defined based on the score at a single blur level $\gamma$, whereas the IEM (and its associated Riemannian metric) integrates over a range of blur levels $\gamma\in[0,{\Gamma}]$. Third, the IEM is constructed from first principles and satisfies several important properties, e.g., it reduces to the well-known Mahalanobis distance when the prior density is Gaussian.
@@
Information geometry (Amari and Nagaoka, 2000) uses tools from differential geometry to analyze statistical models. In particular, it considers probability distributions as points lying on a Riemannian manifold whose metric is derived from the KL divergence. In the case of a family of conditional distributions $\{p_{{\mathbf{y}}|{\mathbf{x}}}(\cdot|{\bm{x}})\,|\,{\bm{x}}\in\mathbb{R}^{d}\}$, this metric is given by the Fisher information matrix (Fisher, 1922), which quantifies the amount of information that $\mathbf{y}$} carries about $\mathbf{x}$} (here, $\mathbf{x}$} is considered as an unknown “parameter”). Specifically, the Fisher information matrix is defined as
@@
Solving a Bayesian inference problem requires both a likelihood function $p_{{\mathbf{y}}|{\mathbf{x}}}$ and a prior $p_{{\mathbf{x}}}$. However, in some cases only the likelihood is available. In such situations, it is natural to choose a non-informative prior using Jeffreys rule (Jeffreys, 1946), which states that the prior should be proportional to the square root of the determinant of the Fisher information matrix. More relevant to our case, this relationship has also been applied in the reverse direction, where the prior is known but the likelihood is not. In particular, it has been shown that a likelihood for which the Jeffreys prior matches the data distribution satisfies the principles of efficient coding (Ganguli and Simoncelli, 2014). The reverse use of Jeffreys prior has also been explored in machine learning. For example, Lebanon (2002) considered a Riemannian metric under which the data is uniformly distributed. Since the prior density is a scalar function, they assumed an isotropic metric of the form $\smash{{\bm{M}}({\bm{x}})\propto\lambda({\bm{x}}){\bm{I}}}$, where $\smash{\lambda({\bm{x}})\propto p({\bm{x}})^{2/d}}$, and ${\bm{x}}\in\mathbb{R}^{d}$. In both of these cases, the result depends only on a scalar quantity and therefore cannot account for the varying magnitudes of discrimination thresholds in different perturbation directions (Berardino et al., 2017). In contrast, the IEM builds on the local geometry of the data distribution to define a distance that captures its anisotropic structure.
@@ A.2 Origins of the pointwise I-MMSE formula
This reformulation highlights that the two sides of the I-MMSE formula correspond to random variables that are equal in expectation. Venkat and Weissman (2012) showed that these random variables satisfy the pointwise I-MMSE formula
@@
where this equality holds with probability one (i.e., almost surely). As noted by Venkat and Weissman (2012), taking expectations in Eq. 2 immediately recovers the original I–MMSE formula. Indeed, the stochastic integral on the right-hand side of Eq. 2 is a martingale with zero mean, while the left-hand side corresponds to the pointwise mutual information between $\mathbf{x}$} and ${\mathbf{y}}_{\gamma}$, whose expectation yields the mutual information between these two random vectors. Using the fact that
@@ Appendix B Mismatched IEM and local decomposition of Kullback–Leibler divergence
The mismatched IEM distance is directly related to the KL divergence between $p_{{\mathbf{x}}_{1}}$ and $p_{{\mathbf{x}}_{2}}$. Indeed, Venkat and Weissman (2012) showed that this KL divergence can be expressed as the expected quadratic variation of the log-probability ratio when the two distributions are evaluated at the same point, $\log\mathopen{}\mathclose{{\left(p_{{\mathbf{y}}_{1,\gamma}}(\gamma{\mathbf{x}}_{1}+{\mathbf{w}}_{\gamma})/p_{{\mathbf{y}}_{2,\gamma}}(\gamma{\mathbf{x}}_{1}+{\mathbf{w}}_{\gamma})}}\right)$, yielding
@@ C.2.2 From the Hessian of the noisy channel log-density to the posterior covariance
This relationship has already appeared in the literature in several contexts, and is often referred to as the “second-order Tweedie identity.” To the best of our knowledge, it was first derived by Hatsell and Nolte (1971, Proposition 3). For completeness and notational consistency, we include a derivation here.
@@ C.2.3 Average local metric
${\bm{e}}_{\gamma}({\bm{x}},{\mathbf{w}}_{\gamma})\coloneqq{\bm{x}}-\mathbb{E}[{\mathbf{x}}|{\mathbf{y}}_{\gamma}=\gamma{\bm{x}}+{\mathbf{w}}_{\gamma}]=-\frac{1}{\gamma}{\mathbf{w}}_{\gamma}-\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma}),$ (45)
@@
$-{\bm{e}}_{\gamma}({\bm{x}}+\bm{\epsilon},{\mathbf{w}}_{\gamma})$ $=\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})+\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\bm{\epsilon}$
$\quad{}+\frac{\gamma^{2}}{2}\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})(\bm{\epsilon},\bm{\epsilon})+o(\mathopen{}\mathclose{{\left\lVert\bm{\epsilon}}}\right\rVert^{2}),$ (46)
@@
$=\frac{1}{2}\int_{0}^{\infty}\mathbb{E}\mathopen{}\mathclose{{\left[2\mathopen{}\mathclose{{\left\langle\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma}),\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\bm{\epsilon}+\frac{\gamma^{2}}{2}\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})(\bm{\epsilon},\bm{\epsilon})}}\right\rangle}}\right$.
$\quad\quad\quad\quad\quad\quad\mathopen{}\mathclose{{\left.+\mathopen{}\mathclose{{\left\lVert\gamma\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\bm{\epsilon}}}\right\rVert^{2}}}\right]\mathrm{d}\gamma+o(\mathopen{}\mathclose{{\left\lVert\bm{\epsilon}}}\right\rVert^{2})$ (47)
$=\bm{\epsilon}^{\top}\int_{0}^{\infty}\mathbb{E}\mathopen{}\mathclose{{\left[\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}}\right)}}\right]\gamma\mathrm{d}\gamma$
$\quad+\frac{1}{2}\bm{\epsilon}^{\top}\mathopen{}\mathclose{{\left(\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma}}\right)\bm{\epsilon}$ (48)
@@
$-\nabla^{2}\log{p_{\mathbf{x}}({\bm{x}})}$ $=\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}(\gamma{\bm{x}}+{\mathbf{w}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma.$ (49)
Taking the expectation over $p_{\mathbf{x}}$, we obtain
$\mathbb{E}\mathopen{}\mathclose{{\left[-\nabla^{2}\log p_{\mathbf{x}}({\mathbf{x}})}}\right]$ $=\int^{\infty}_{0}\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(\nabla^{2}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)^{2}+\nabla^{3}\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})\mathopen{}\mathclose{{\left(\frac{1}{\gamma}{\mathbf{w}}_{\gamma}+\nabla\log p_{{\mathbf{y}}_{\gamma}}({\mathbf{y}}_{\gamma})}}\right)}}\right]\gamma^{2}\mathrm{d}\gamma,$ (50)
where we used the fact that ${\mathbf{y}}_{\gamma}=\gamma{\mathbf{x}}+{\mathbf{w}}_{\gamma}$. The second term in the expectation then vanishes due to Stein’s lemma (Stein, 1981),
@@
The last equality in the theorem is generic and classical. By expanding the derivatives, we have
$-\int p_{{\mathbf{x}}}({\bm{x}})\nabla^{2}\log p_{{\mathbf{x}}}({\bm{x}})\mathrm{d}{\bm{x}}$ $=\int\mathopen{}\mathclose{{\left(p_{{\mathbf{x}}}({\bm{x}})\nabla\log p_{\mathbf{x}}({\bm{x}})\nabla\log p_{\mathbf{x}}({\bm{x}})^{\top}-\nabla^{2}p_{\mathbf{x}}({\bm{x}})}}\right)\mathrm{d}{\bm{x}}.$ (54)
The second term in the integral on the right-hand side vanishes through integration by parts.